The slope of the line passing through P(1,4) and Q(x,1) is -3/5. What is the value of x?
a. 6
b. 7
c. 9
d. 10

Answer by Timnewman(323)

Recall that the slope of a line is given by:
y2-y1=m(x2-x1)
but,y2=1,y1=4,x2=x,x1=1,m=-3/5
substitute the above in the formula,
1-4=-3/5(x-1)
-3=(-3x+3)/5
multiply bothsides by 5
-15=-3x+3
3x=3+15
3x=18
x=18/3
x=6
The value of x is 6
therefore,option (a) is the key.
HOPE THIS HELPS,,,,,,
----TIMNEWMAN